### v1

# references.bib

references.bib
Solid state quantum emitters are a prime candidate in distributed quantum technologies since they inherently provide a spin-photon interface. An ongoing challenge in the field, however, is the low photon extraction due to the high refractive index of typical host materials. This challenge can be overcome using photonic structures. Here, we report the integration of V2 centers in a cavity-based optical antenna. The structure consists of a silver-coated, -thin 4H-SiC membrane functioning as a planar cavity with a broadband resonance yielding a theoretical photon collection enhancement factor of 34. The planar geometry allows us to identify over 20 single V2 centers at room temperature with a mean (maximum) count rate enhancement factor of 9 (15). Moreover, we observe 10 V2 centers with a mean absorption linewidth below at cryogenic temperatures. These results demonstrate a photon collection enhancement that is robust to the lateral emitter position.
INTRODUCTION
Optically active spins in solid-state materials are promising candidates for applications in quantum technologies [Awschalom2013, Awschalom2018, Atatuere2018]. Besides the well-established diamond platform [Jelezko2006, Doherty2013], 4H-silicon carbide (SiC) has gained growing attention in the field. It hosts several promising emitters such as the silicon-vacancy center (V) [Kraus2014, Widmann2014, Nagy2019], the divacancy center (VV) [Son2006, Christle2017, Li2022], as well as the more recently investigated nitrogen-vacancy center [VonBardeleben2015, Wang2020] and the vanadium center [Wolfowicz2020, Cilibrizzi2023]. This has yielded a number of encouraging results with regards to quantum networks, such as single shot readout of the electron spin state, through charge state conversion using the VV center [Anderson2022], or through a nearby nuclear spin in case of the V center [Lai2024, Hesselmeier2024], as well as spin-photon entanglement using the V [Fang2024].
In addition, silicon carbide offers excellent nanofabrication capabilities. This enables implementing various photonic structures that improve the rate of collected photons from emitters through their geometry, e.g., solid-immersion lenses (SIL) [Sardi2020, Bekker2023], nanopillars [Radulaski2017], and waveguides [Babin2021, Krumrein2024]. It also allows for the fabrication of more complex structures such as photonic crystal cavities [Bracher2017, Lukin2020], or disk resonators [Wang2021, Lukin2023] that can further boost the photon rate through Purcell enhancement. While nanophotonic cavities offer the highest enhancement among the above mentioned structures, they rely on complicated fabrication processes, tight restrictions on the emitter placement, and complex methods to keep the emitter on resonance with the structure.
Approaches like bullseye cavities [Li2015, Hekmati2023], plasmonic enhancement structures [Shalaginov2019, Zhou2023], and planar antenna structures [Lee2011, Checcucci2017, Fuchs2021] represent an interesting middle ground. They have shown a much higher photon count enhancement than SILs or nanopillars while offering relaxed restrictions on the emitter position. Additionally, the operation conditions of such structures are less complicated since they work over a broader spectral range at the cost of a smaller count rate enhancement in comparison to nanophotonic cavities.
Inspired by previous results on tin vacancies in a diamond structure [Fuchs2021], we demonstrate the successful integration of single V2 centers (V centers at the cubic lattice site in 4H-SiC) into a cavity-based optical antenna, consisting of a -thin 4H-SiC membrane with silver coatings on both sides. We investigate 25 single V2 centers in a large region of the antenna () at room temperature and demonstrate an average fluorescence enhancement of factor 9 with a maximum of factor 15. Additionally, we show that the ODMR contrast of V2 centers in the structure is similar to the one of emitters in a bulk sample. Further, we investigate 10 V2 centers in the antenna at cryogenic temperatures below and show resonant absorption of the two optical transitions with mean linewidths below and an inhomogeneous ZPL distribution of , which is comparable to our previous investigation of V2 centers in membranes [Heiler2024]. Lastly, we investigate the photon collection enhancement at low temperatures and find an increase in the photon count rate by a factor of 6.6 for the brightest emitter.
Compared to SILs or nanopillars, our structure offers a fluorescence enhancement over a much larger spatial region, only limited by the homogeneity of the membrane thickness. Our results provide a promising route to strongly enhanced fluorescence with spin-selective excitation and thus could be used to boost readout rates in electron-nuclear spin experiments [Lai2024, Hesselmeier2024]. As shown in [Fuchs2021], our design could be adapted to a smaller resonance around the ZPL with an increased membrane thickness, i.e. with expectedly smaller absorption linewidths [Heiler2024], at the cost of a slightly reduced enhancement. Thus, it is also of high interest for applications in quantum networking, where a high ZPL rate and coherent optical lines are key requirements [Northup2014, Ruf2021].
RESULTS AND DISCUSSION
Cavity antenna optimization and fabrication
A schematic of the cavity antenna is shown in Figure 1(a). The structure is based on a sub-micron-thin 4H-SiC membrane with silver coatings on both sides. Due to the high reflectance of the silver, the structure acts as a Fabry-Pérot cavity. Choosing the thickness of the upper silver layer well below makes it transparent enough such that the cavity modes can be converted to free-propagating modes in the upper direction. The lower silver layer is kept thick enough (several ) to fully reflect the light. As already discussed in [Fuchs2021], this results in a higher photon emission rate and an increased directivity towards the collectable solid angle of an objective for the electric dipole radiation of a color center inside the SiC membrane. Finally, the upper silver layer is capped by an SiO layer to prevent fast oxidation of the silver at ambient conditions. Its thickness is optimized to work as an anti-reflective coating [Fuchs2021]. Note, that plasmonic resonances at the interfaces are a loss channel for this structure rather than the enhancement mechanism like in other approaches [Shalaginov2019, Zhou2023].
In contrast to the work of [Fuchs2021], the aim of our structure is to maximize the collectable photon rate within the phonon-side band (PSB) of the color center, in our case the V2 center. Therefore, we optimize the structure to achieve a very broad resonance in the wavelength range of its PSB without including a second resonance that could be used for a more efficient off-resonant excitation. This is enabled by tuning the membrane thickness to the first fundamental mode of the cavity, as depicted in Figure 1(e).
For a given set of antenna layer thicknesses, the wavelength dependent emission enhancement over an emitter in bulk crystal is shown as the orange curve in Figure 1(b). We numerically optimize the thickness of the upper silver layer, the membrane, the SiO capping layer, and the distance of the dipole from the lower silver layer to yield a maximum total enhancement [Software2021] (see Supporting Information for details). The total enhancement is calculated by weighting the wavelength dependent enhancement over the normalized fluorescence spectrum of the V2 center, shown in blue in Figure 1(b). Figure 1(c) shows the total enhancement for the collection within a spectral band of as a function of the two most relevant thicknesses in the design, i.e., the upper silver layer and the membrane thickness. We choose this specific wavelength range, since our detector becomes much less efficient for wavelengths above (see Supporting Information). The structure yields a significant enhancement for membrane thicknesses between and upper silver thicknesses between with a maximum enhancement factor of 34. To study the influence of the dipole position within the thin SiC membrane, we simulate the enhancement factor as a function of the membrane thickness and the relative dipole position at the optimized upper silver layer thickness of in Figure 1(e). For the first cavity mode, the enhancement is very robust to a displacement of the dipole from its optimal position in the middle of the membrane. It stays in between for a relative dipole position in the membrane in between from the lower silver-SiC interface. The robustness of the enhancement on more parameters is shown in the Supporting Information.
The fabrication of the antenna structure starts with the steps described in our previous work [Heiler2024] in order to create sub-µm thin 4H-SiC membranes with integrated V2-centers. With this, we create a sample with a total thickness of containing several subregions with a thickness below , which we refer to as membranes in the following. In the last step of this process, we etch our sample down to a thickness of in the central part of one membrane, measured by a home-built white light interferometer (see Supporting Information for details on the membrane thickness). The membrane is subsequently coated using electron beam evaporation with a target silver thickness of for the thick silver layer (full-reflective mirror) and for the thin silver layer (leaky mirror). In the final step, the SiO layer is sputtered on top of the thin silver layer. To check the resonance of the fabricated structure, we measure its reflectivity spectrum after the coating steps using the WLI. The result, normalised by a reference spectrum from a mirror, is depicted in Figure 1(d) and shows a reflection dip at that covers a big part of the phonon side band of the V2 center. We can model the measured data using a transfer matrix model and find a good agreement for an upper silver layer of and a SiC thickness of , confirming the earlier WLI measurement.
Fluorescence enhancement at room temperature
To test the enhancement of V2 fluorescence in our structure, we investigate the sample in a confocal laser scanning setup at room temperature (details in the Supporting Information). At first, we measure fluorescence from the central part of the antenna and find a region with a diameter of around the membrane center that shows bright fluorescence, depicted as the black circle in Figure 2(a). The diameter of the enhanced region is determined by the thickness variation of the SiC membrane in this region. We measured this variation to be (see Supporting Information for details), shown as white dashed lines in Figures 1(c) and 1(e), and see the corresponding enhancement drop outside of this region. A zoom of the region reveals the presence of diffraction limited spots that appear significantly brighter than the spots in a scan taken on a bulk reference sample at similar excitation conditions, see Figure 2(b). To be comparable, V2 centers in the reference sample were created with electron irradiation at the same dose as for the antenna sample.
To confirm the presence of single V2 centers in the antenna sample, we continue with optically-detected magnetic resonance (ODMR) and autocorrelation measurements on the small, bright spots we find in a confocal scan of the sample (experimental details can be found in the Supporting Information). Figures 2(c) and 2(d) show the results of these measurements on an exemplary V2 center in our structure. We find an ODMR contrast of , which is a typical value for the V2 center at room temperature [Radulaski2017, Wang2017, Singh2023, Krumrein2024] and an anti-bunching of g, significantly below at zero time delay, without background correction, confirming that we are investigating a single V2-center. We note that the typically observed single ODMR peak for the V2-center without an applied magnetic field is split by into two peaks in our case. Since this behavior is found for most V2-centers in different samples investigated on the setup with splittings between (see Supporting Information), we attribute it to a magnetization of parts of our setup.
To quantify the enhancement of the photon count rate of the exemplary emitter, we perform excitation power dependent measurements of the count rate on it and fit the saturation count rate to . The same measurements are performed on a single V2 center in a reference bulk sample and on a single emitter in a SIL (see Supporting Information for details on the reference samples), yielding saturation counts of and , as shown in Figure 2(e). These measurements show that the emitter in the antenna is enhanced by more than one order of magnitude compared to the bulk emitter and a factor of compared to the SIL emitter.
For better statistics on the enhancement factor, we investigate a higher number of emitters. At first, we record ODMR for individual spots on the bulk and antenna sample to find V2 centers. We set a threshold of at least contrast in the ODMR to use an emitter for further measurements, since a much lower contrast compared to the average indicates fluorescence from different defects, e.g., the V1 center, or crystal damage in the spot. With this, we find that of the bright spots in the antenna are V2 centers which is lower compared to in the bulk reference sample. We suspect the close proximity to the surface in our antenna sample yields a higher number of surface related defects in confocal scans, thus, decreases the relative number of V2 emitters. However, with a mean density of we still find a high number of V2 centers in the antenna sample. For a faster identification of V2 centers among the bright spots, we use a pre-selection based on the polarization of the emission (see Supporting Information for details). In the following, we use 25 representative emitters in the antenna, 23 in the bulk and 7 emitters in the SIL sample.
To ensure that only single emitters are included into our statistic, we measure the second order autocorrelation function. For the 25 antenna emitters, we find 14 (21) emitters with an autocorrelation dip below 0.5 before (after) applying background correction. Due to the low brightness, we only test 7 of the 23 emitters in bulk for autocorrelation and all of them show a clear anti-bunching below 0.5 after background correction. The 7 emitters from the SIL sample show anti-bunching below 0.5 after correction for the background as well. The autocorrelation values for all emitters of the study can be found in the Supporting Information.
Subsequently, we measure excitation power dependent photon count rates for all single V2 centers to extract the saturation counts as described in the Supporting Information. Figure 2(f) shows a histogram of the saturation count rates for all emitters from the bulk, SIL, and antenna sample. Including all emitters with an anti-bunching below 0.5 after applying background correction, we find an average enhancement of () for the antenna sample which is significantly higher than the average enhancement of () from the SIL sample. The maximum enhancement is ( for the emitter with the highest saturation count rate in the antenna.
Absorption linewidths and inhomogeneous distribution of ZPLs under resonant excitation
Since applications based on a spin–photon interface require an operation under cryogenic conditions, we investigate the optical properties of antenna-integrated V2 centers at temperatures below . To identify the V2 centers within diffraction limited spots, we measure their emission spectrum at off-resonant excitation in a low temperature confocal microscope setup (details see Supporting Information). We investigate 32 representative emitters with a distinct emission peak in the range of by photoluminescence excitation (PLE) measurements. Of those emitters, show two clearly distinguishable optical spin-conserving transitions A and A, as expected for the V2 center [Kraus2014], while show a broad PLE and no PLE at all (see Supporting Information for details).
We note that we were able to measure PLE without an external microwave drive or a static magnetic field, which are typically necessary to counteract optical spin pumping [Banks2019, Babin2021]. This indicates a spin-mixing of the ground state sub-levels which we suspect to originate from an increased amount of strain at low temperatures due to different thermal expansion of the layers.
At resonant excitation with low power of (), we find emitters with a stable PLE over a long time scale, as shown for a representative emitter in 3(a) for half an hour. We perform excitation power dependent PLE measurements to estimate the absorption linewidths of 10 arbitrarily selected emitters out of the emitters with distinguishable transitions. Each emitter is measured for 10 single sweeps at up to six different excitation powers in between . The mean A and A linewidths for each power obtained by averaging the linewidth fit parameter from a double peak Gaussian fit of each single sweep are depicted in 3(b). For comparison, the mean A and A linewidths of an emitter in a SIL sample are measured as well, using a spin-mixing microwave signal. As expected due to the much larger distance of the emitter from the SiC surface, the linewidths from the SIL emitter lie below the respective antenna sample transition linewidths, except for the highest excitation power, where power broadening becomes significant for both structures. In total, the average linewidth for each power for the antenna emitters lies below . Data on the power dependent linewidths of each individual antenna emitter can be found in the Supporting Information.
The mean spectral position of the A peak for each measured emitter in the antenna compared to the respective position of the emitter in the SIL is depicted in Figure 3(c). The relative positions range from to , spanning a total of . A smaller range of was already observed in membranes [Heiler2024]. For comparison, the spectral range of A peak positions in 40 measured SIL emitters from another sample, also shown in Figure 3(c), amounts to . The bigger spread in thinner membrane structures can be explained by the increased influence of fabrication and strain on the integrated emitters.
Next, we quantify the spectral wandering as the positional change of the A peak in consecutive single sweeps, to get a measure of the spectral stability over time. This data is presented in form of histograms in Figure 3(d). As expected, the standard deviation of spectral wandering of emitters in the membrane with is much bigger compared to the SIL emitter with where the emitter is basically in bulk crystal environment. This observation also aligns with previous work [Heiler2024].
Finally, we investigate the count rate enhancement of the structure under resonant excitation. To prevent a (partial) spin-polarization, which would lower the amount of collected photons, we scan two lasers separated by simultaneously to drive the A and A transition at the same time (details in the Supporting Information). With this excitation scheme, we record power-dependent PLE for each three emitters in the reference bulk and SIL sample and 10 emitters in the antenna sample up to an excitation power where the emitter ionizes already after a single PLE scan most of the time. We evaluate the photon count rate for each excitation power by averaging the amplitudes from Gaussian fits on the single PLE lines. However, we find that especially for the emitters in the antenna sample, the power dependent average count rates do not reach the typical non-linear regime of saturation since they ionize in almost every line before. This effect results in saturation curves where no saturation power can be confidently extracted via a fit apart from 4 antenna emitters (see Supporting Information).
The extracted average enhancement is for the SIL and for the antenna sample. To get an estimate of the maximum enhancement, we compare the brightest single PLE line with no sign of ionization from the bulk, SIL and antenna emitters in Figure 3(e). We extract an enhancement of for the brightest SIL and for the brightest antenna emitter. The average enhancement in the antenna at low temperatures is of the enhancement extracted at room temperature. This decrease is expected due to the higher detection efficiency at the low temperature setup for wavelengths above where the structure shows a smaller enhancement (see Supporting Information for details).
Summary and Outlook
In summary, we optimized and fabricated a cavity-based 4H-SiC antenna structure and demonstrated a collection count enhancement of factor 9 (15) in average (at most) at room temperature from integrated V2 centers. At low temperatures, we found an average absorption linewidth below and a collection count enhancement of factor 6.6 for the brightest checked emitter. Further, we find a standard deviation in the spectral position over time of which remains small enough to implement a resonance feedback system reported previously [Babin2021].
Compared to nanopillars, SILs, or nanobeam resonators that enhance emitters in a small spatial region, our structure offers the big advantage to enhance emitters in a two-dimensional plane, that is limited in size solely by the fabrication. Thus, further work on the fabrication, i.e. creating all color centers in the optimal depth using ion-implantation methods [Ohshima2018, Wang2019] and further optimizing the thickness homogeneity, can be used to significantly scale up the number of enhanced emitters in a single structure. This simplifies the integration of defects with a low creation yield, e.g., SiC divacancies near stacking faults [He2024]. Moreover, if the structure is transferred to a system with emitters that have a broad inhomogeneous linewidths, e.g., optically active molecules [Checcucci2017, Toninelli2021] or rare earth ion quantum emitters [Thiel2011], the big number of enhanced emitters eases finding emitters with matching emission lines, enabling two-emitter entanglement. The versatile design also enables enhancing the ZPL using the spectrally smaller resonance of a thicker membrane and making the structure very appealing in the context of quantum communication where indistinguishable photons are key.
DATA AVAILABILITY
The data supporting the presented findings are available from the corresponding author upon reasonable request.
CODE AVAILABILITY
The measurement and evaluation codes used for this study are available from the corresponding author upon reasonable request. The optimization of our optical structure is based on the published python code [Software2021].
AUTHOR CONTRIBUTIONS
J.K. and J.H. contributed equally to this work. The project was conceived by J.K., J.H., P.Fu., C.B., and F.K. and supervised by F.K. and J.W. The high-quality epilayer was grown by J.U.H. and the sample was electron irradiated by W.K. J.K. and J.H. fabricated the SiC membrane and performed fabrication related measurements of the sample. P.Fl. applied the thin-silver layer and the SiO coating. The optical measurements at room temperatures were conducted and analyzed by J.K. The optical measurements at low temperatures were conducted by J.K., E.H., and P.K. and analyzed by J.H. and J.K. The reference sample was fabricated by E.H. The manuscript was written by J.K. and J.H. All authors contributed to the manuscript.
COMPETING INTERESTS
The authors declare no competing interests.
ACKNOWLEDGEMENTS
We acknowledge fruitful discussions and experimental help from M. Krumrein, R. Stöhr, and A. Weible.
J.H., F.K., and J.W. acknowledge support from the European Commission for the Quantum Technology Flagship project QIA (Grant agreements No. 101080128, and 101102140). P.K., F.K., J.U.H, and J.W. acknowledge support from the European Commission through the QuantERA project InQuRe (Grant agreements No. 731473, and 101017733). P.K., F.K. and J.W. acknowledge the German ministry of education and research for the project InQuRe (BMBF, Grant agreement No. 16KIS1639K). C.B., F.K. and J.W. further acknowledge the German ministry of education and research for the project QR.X (BMBF, Grant agreements No. 16KISQ001K and 16KISQ013), while J.W. also acknowledges support from the EU Horizon project SPINUS (project: 101135699). F.K. and J.W. additionally acknowledge the Baden-Württemberg Stiftung for the project SPOC (Grant agreement No. QT-6). F.K. acknowledges funding by the Luxembourg National Research Fund (FNR, project: 17792569) in addition. J.U.H further acknowledges support from the Swedish Research Council under VR Grant No. 2020-05444 and Knut and Alice Wallenberg Foundation (Grant No. KAW 2018.0071).